Indiana was hired to investigate a temple ruin on a desolated island within the Bermuda triangle. He flew a seaplane going  north towards the island. As he approached the island, the flight control tower notified him that an unauthorized plane is flying  south at about the same altitude and located on the same area. Indiana was informed that the southward bound plane is  currently located at a position which is 13.5 kilometers, 102˚ from his plane.  


a. How far to the north is the other plane? 

b. How far to the west is the other plane? 

c. Afraid that the plane was an enemy aircraft, Indiana asked how long will it take before their planes meets side-by-side.  The control tower responded that the other plane is traveling at the same speed as Indiana at 290 km/hr. Suddenly,  Indiana lost connection with the control tower. Help Indiana to determine how much time will elapse before their planes  are side-by-side. 




1
Expert's answer
2021-02-17T12:06:12-0500

a)


"s_y=-13.5\cos{102}=2.8\ km"

b)


"s_x=13.5\sin{102}=13.2\ km"

c)


"t=\frac{13.5}{290}60=2.8\ min"